# Pivoting technique for the circle homeomorphism group

### v1

@@ 4. Schottky sets and random walks
$\zeta(S)$ $:=\sup\big{\{}\zeta(I_{1}),\ldots,\zeta(I_{N}),\zeta(J_{1}),\ldots,\zeta(J_{N}),\big{\}}$
@@
#$\#\big{(}\mathcal{A}:=\{i:\textrm{$I_{i}$ contains $x$ or $y$}\}\big{)}\leq 2.$
@@
Hence, $\mathcal{A}\cup\mathcal{B}$ has at most #$\#\sqrt{N}+2<N$ elements, and we can pick an index $i$ outside it. For that $i$, we conclude that $d(gx,gy)<\operatorname{Leb}(J_{i})\leq 1/\sqrt{N}$ for each $g\in\mathfrak{S}(I_{i},J_{i})$. Since $G$ intersects with $\mathfrak{S}(I_{i},J_{i})$, we conclude that $\inf_{g\in G}d(gx,gy)<1/\sqrt{N}$. By sending $N$ to infinity, we conclude that the action of $G$ is proximal. ∎
@@ 4.1. Exponential Synchronization
First, let us write $S=\mathfrak{S}(I_{1},J_{1})\cup\ldots\cup\mathfrak{S}(I_{N},J_{N})$ for some essentially disjoint sets $I_{1},\ldots,I_{N},J_{1},\ldots,J_{N}$. Recall that elements of $\mathfrak{S}(I_{i},J_{i})$ send $\mathcal{I}$ into $J_{i}$. ()($\ast$)( ∗ ) Note that $wJ_{1},\ldots,w_{J_{N}}$ are disjointly contained in $w\mathcal{I}$. Hence, the sum of their “size” is no greater than that of $w\mathcal{I}$, which implies that
@@
Now fixing $10\leq k\leq n-$ 1 and the choices of $\{s_{i}:1\leq i\leq k\}$, we observe that
@@
($\operatorname{\mathbb{P}}\big{(}\{x,y\}\cap w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\mathcal{I}=\emptyset\big{)}\geq 1-e^{-n}$
@@
holds regardless of the choices of $s_{i}$’s. Furthermore, when $10\leq k\leq n-$ 1 and $\{s_{i}:1\leq i\leq k\}$ are given,
@@
($\operatorname{\mathbb{P}}\big{(}\{x,y\}\cap W_{n}\mathcal{I}=\emptyset\big{)}\geq\operatorname{\mathbb{P}}\left(E_{1}\cup\ldots\cup E_{n}\right)\geq 1-(2/N)^{n}\geq 1-e^{-n}.$
@@ 4.2. Probabilistic Tits alternative
#$\#\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$ intersect}\big{\}}\leq 3N\sqrt{N}.$
@@
#$\geq\#\big{\{}(I_{i},gJ_{j}):I_{i}\cap gJ_{j}\neq\emptyset,I_{i}\in\mathcal{C}\big{\}}$
#$\geq\#\mathcal{C}\cdot\min_{I_{i}\in\mathcal{C}}\#\{J_{j}:I_{i}\cap gJ_{j}\neq\emptyset\}$
#$\geq\#\mathcal{C}\sqrt{N}$
@@
#$\#\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$intersect}\big{\}}$$\leq\#\mathcal{C}\cdot N+(N-\#\mathcal{C})\cdot\sqrt{N}$
@@
#$\#\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$ intersect}\big{\}}\leq 3\zeta^{2}N\sqrt{N}.$
@@
Hence, if $i\notin\mathcal{C}^{(l,m)}$ for each $(l,m)\in\{1,\ldots,\zeta\}^{2}$, then $\{j:I_{i}\cap gJ_{j}\neq\emptyset\}$ has cardinality $\leq\zeta^{2}\sqrt{N}$. Since #$\#\mathcal{C}^{(l,m)}$ has cardinality at most $2\sqrt{N}$ for each $(l,m)$, we conclude that $\mathcal{C}$ consists of at most $2\zeta^{2}\sqrt{N}$ sets among $\{I_{1},\ldots,I_{N}\}$.
@@
#$\#\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$intersect}\big{\}}$$\leq\#\mathcal{C}\cdot N+(N-\#\mathcal{C})\cdot\zeta^{2}\sqrt{N}$
@@
(w0s1w1s2nw2n and v0t1v1t2nv2n comprisea Schottky pair and generate a free $\operatorname{\mathbb{P}}\left(\begin{array}[]{c}\textrm{$w_{0}s_{1}w_{1}\cdots s_{2n}w_{2n}$ and $v_{0}t_{1}v_{1}\cdots t_{2n}v_{2n}$ comprise}\\ \textrm{a Schottky pair and generate a free subgroup}\end{array}\right)\geq 1-6e^{-n/10}.$
@@
$E_{1}$ :=$:=\big{\{}s_{n+1}w_{n+1}\cdots s_{2n}w_{2n}\cdot w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\bar{\mathcal{I}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}\big{\}},$
$E_{2}$ :=$:=\big{\{}t_{n+1}v_{n+1}\cdots t_{2n}v_{2n}\cdot v_{0}t_{1}v_{1}\cdots t_{n}v_{n}\bar{\mathcal{I}^{\prime}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}^{\prime}\big{\}},$
$E_{3}$ :=$:=\big{\{}s_{n+1}w_{n+1}\cdots s_{2n}w_{2n}\cdot v_{0}t_{1}v_{1}\cdots t_{n}v_{n}\bar{\mathcal{I}^{\prime}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}\big{\}},$
$E_{4}$ :=$:=\big{\{}t_{n+1}v_{n+1}\cdots t_{2n}v_{2n}\cdot w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\bar{\mathcal{I}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}^{\prime}\big{\}},$
$E_{5}$ :=$:=\big{\{}s_{n+1}w_{n+1}\cdots s_{2n}w_{2n}\cdot v_{2n}^{-1}t_{2n}^{-1}\cdots v_{n+1}^{-1}t_{n+1}^{-1}\cdot\overline{S^{1}\setminus\mathcal{I}^{\prime}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}\big{\}},$
$E_{6}$ :=$:=\big{\{}v_{n}^{-1}t_{n}^{-1}\cdots v_{1}^{-1}t_{1}^{-1}v_{0}^{-1}\cdot w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\cdot\bar{\mathcal{I}}\subseteq\operatorname{\operatorname{int}}(S^{1}\setminus\mathcal{I}^{\prime})\big{\}}.$
@@
$I^{(1)}$ $:=(s_{n+1}w_{n+1}\cdots s_{2n}w_{2n})^{-1}(S^{1}\setminus\operatorname{\operatorname{int}}\mathcal{I}),$
$I^{(2)}$ $:=(t_{n+1}v_{n+1}\cdots t_{2n}v_{2n})^{-1}\cdot(S^{1}\setminus\operatorname{\operatorname{int}}\mathcal{I}^{\prime}),$
$J^{(1)}$ $:=w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\overline{\mathcal{I}},$
$J^{(2)}$ $:=v_{0}t_{1}v_{1}\cdots t_{n}v_{n}\overline{\mathcal{I}^{\prime}}.$
@@
$\operatorname{\mathbb{P}}_{(Z_{n},Z_{n}^{\prime})\sim\mu^{\ast n}\times\mu^{\prime\ast n}}\big{(}\textrm{$Z_{n},Z_{n}^{\prime}$ comprise a Schottky pair and generate a free subgroup}\big{)}\geq 1-\frac{1}{\kappa_{2}}e^{-\kappa_{2}n}.$
@@ 4.3. Local contraction
×$\cdots\times\Omega_{3}\times\Omega_{2}\times\Omega_{1}=:\Omega,$
@@ 5. Pivoting technique
It suffices to prove this for $n$ being a multiple of $3m$. Indeed, for $n=3mk+l$ ($11\leq l\leq 3m-$ 1) we can treat as follows: we first take $\Omega_{3mk}$, $\mathcal{P}_{mk}$, $(w_{i})_{i},(r_{i},s_{i},t_{i})_{i}$ using the proposition and consider $(G^{l},\mu^{l})$ (where $G=\operatorname{Homeo}(S^{1})$). And then we define
$\Omega_{3mk+l}$ $:=\Omega_{mk}\times G^{l},$
$\mathcal{P}_{3mk+l}$ $:=\mathcal{P}_{mk}\times G^{l}=\{\mathcal{E}_{\alpha}\times(g_{1},\ldots,g_{l}):\mathcal{E}_{\alpha}\in\mathcal{P}_{3mk},(g_{1},\ldots,g_{l})\in G^{l}\}$
@@
$w_{0}$ $:=\prod_{i=1}^{i(1)-1}g_{i}=\nu_{1}\cdots\nu_{i(1)-1},$
$w_{l}$ $:=\prod_{i=i(l)+1}^{i(l+1)-1}g_{i}=\nu_{i(l)+1}\cdots\nu_{i(l+1)-1}$ $(l=1,\ldots,\lfloor\epsilon n/10m\rfloor-1)$
$w_{\lfloor\epsilon n/10m\rfloor}$ $:=\prod_{i=i(\lfloor\epsilon n/10m\rfloor)+1}^{n/3m}g_{i}=\nu_{i(\lfloor\epsilon n/10m\rfloor)+1}\cdots\nu_{n/3m}$
@@
First, note that the elements of $P_{u}$ smaller than or equal to $m-1$ must have been acquired no later than step $m-1$, and then must have never been lost thereafter. Hence, they all belong to $P_{m-1}$. Meanwhile, all elements $P_{m-1}$ should have remained till step $u$ for the following reason. If an element of $P_{m-1}\subseteq\{1,\ldots,m-1\}$ was lost at some step $n\in\{m,m+1,\ldots,u\}$, it would mean that scenario (2-B) was the case at step $n$, with $k=\max\mathcal{Q}$ being smaller than $m-1$. This means that $P_{n}$ lost not only $P_{m-1}$ but also $m$, which contradicts $P_{u}\ni m$. Hence the claim follows.
@@
    $l=m-1$: this means that scenario (2-A) was the case at both step $l$ and step $m=l+1$. Hence, $J(r_{l+1})\subseteq W_{l}^{-1}L_{l}:=w_{l}^{-1}\big{(}S^{1}\setminus I(t_{l})\big{)}$ must hold. This implies
@@
    $l<m-1$: in this case, $P_{m-1}=P_{u}\cap\{1,\ldots,m-1\}\subseteq\{1,\ldots,l\}$ does not contain $m-1$ so scenario (2-B) must have been the case at step $n=m-1$. Still, $P_{m-1}=P_{u}\cap\{1,\ldots,m-1\}$ contains an element $l$ so scenario (2-B-ii) is ruled out. Thus, scenario (2-B-i) was the case and $l$ must have been the maximum element of $\mathcal{Q}$. This leads to $L_{m-1}:=W_{l-1}r_{l}s_{l}\mathcal{I}$. We now know that scenario (2-A) was the case at step $n=m$, which implies $J(r_{m})\subseteq W_{m}^{-1}L_{m-1}$. Hence, we conclude
@@
Let $S=\cup_{i=1}^{N}\mathfrak{S}(I_{i},J_{i})$ for essentially disjoint subsets $\{I_{i},J_{i}\}_{i}$. Note that the set $P_{n-1}$ and the interval $L_{n-1}$ are determined from the fixed inputs. Now at step $n-1$ of the pivotal set construction, three possibilities arise:
@@
Based on our estimates for the probabilities for $I(t_{n})\notin\mathcal{C}(w_{n};S)$ and $J(r_{n})\subseteq W_{n-1}^{-1}L_{n-1}$ in the above three cases, we can conclude that #$\#P_{n+1}=\#P_{n}+1$ happens for probability at least $1-4\zeta^{2}/\sqrt{N}$. ∎
@@
[…]
@@
As the base case, we have observed Item (4) for $l=k$. For general $k<l\leq n$, we will start by assuming Item(4) for $l-1$. Recall the conditions for scenario (2-A) at step $l$, for the input $\mathbf{s}$:
@@
$\big{[}(\mathbf{r},\mathbf{s},\mathbf{t})\sim_{n}(\bar{\mathbf{r}},\bar{\mathbf{s}},\bar{\mathbf{t}})\big{]}\Leftrightarrow\left[\begin{array}[]{c}\textrm{$r_{i}=\bar{r}_{i}$ and $t_{i}=\bar{t}_{i}$ for each $i\in\operatorname{\mathbb{Z}}_{>0}\setminus\{n+1\}$ and}\\ \textrm{$\bar{s}_{i}=s_{i}$ for each $i\in\operatorname{\mathbb{Z}}_{>0}\setminus P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})$}\end{array}\right].$
@@
[…]
@@
It remains to prove the claim. The claim regarding $A_{0}$ was already established in Lemma 5.10. That means, regardless of the values of $(s_{i(M)},\ldots,s_{i(1)})$, we proved that the probability for $(r_{n+1},t_{n+1})$ to satisfy #$\#P_{n+1}=\#P_{n}+1$ is at least $1-4\zeta^{2}/\sqrt{N}$. We will prove something more: we claim that the candidates for $r_{n+1},t_{n+1}$ that make #$\#P_{n+1}=\#P_{n}+1$ are independent of $(s_{i(M)},\ldots,s_{i(1)})$. When restricted to $\mathcal{E}$, #$\#P_{n+1}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})=\#P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})+1$ holds if and only if $I(t_{n+1})\in\mathcal{C}(w_{n+1};S)$ and $J(r_{n+1})\subseteq W_{n}^{-1}L_{n}$. Here,
@@
are fixed throughout $\mathcal{E}$. This is why #$\#P_{n+1}=\#P_{n}+1$ depends on the choice of $r_{n+1}$ and $t_{n+1}$, regardless of the values of $s_{i(1)},\ldots,s_{i(M)}$. This settles Claim 5.15(1), (2) and also the construction of $A_{0}$.
@@
$A_{l}$ :=$:=\Big{\{}s\in S:I(s)\notin\mathcal{C}\big{(}t_{i(l)}w_{i(l)}\cdot W_{i(l)}^{-1}W_{n+1};S\big{)}\Big{\}}$
=$=\Big{\{}s\in S:I(s)\notin\mathcal{C}\big{(}t_{i(l)}w_{i(l)}\cdot(r_{i(l)+1}s_{i(l)+1}t_{i(l)+1}w_{i(l)+1})\cdots(r_{n}s_{n}t_{n}w_{n})\cdot(r_{n+1}s_{n+1}t_{n+1}w_{n+1});S\big{)}\Big{\}}.$
@@
(5.2) #$\#\big{\{}j:\bar{I}(s_{i(l)})\cap t_{i(l)}w_{i(l)}\cdot W_{i(l)}^{-1}W_{n+1}\bar{J}_{j}\neq\emptyset\big{\}}\leq\zeta^{2}\sqrt{N}.$
@@
#$\#\big{\{}j:\bar{I}(t_{i(l+1)})\cap(W_{i(l+1)}w_{i(l+1)}^{-1})^{-1}\cdot(W_{i(l)-1}r_{i(l)}s_{i(l)})\cdot t_{i(l)}w_{i(l)}\cdot W_{i(l)}^{-1}W_{n+1}\bar{J}_{j}\neq\emptyset\big{\}}\leq\sqrt{N}.$
In other words, $I(t_{i(l+1)})\notin\mathcal{C}\big{(}w_{i(l+1)}\cdot(W_{i(l+1)})^{-1}\cdot W_{n+1};S\big{)}$ holds true. This implies that the set $\mathcal{Q}$ in scenario (2-B) at step $n+1$ contains $i(l+1)$. Hence, $P_{n+1}(\mathbf{r},\mathbf{s},\mathbf{t})$ contains $P_{n}(\mathcal{E})\cap\{1,\ldots,i(l+1)\}=\{i(M)<\ldots<i(l+1)\}$ at least, which leads to the inequality #$\#P_{n+1}\geq\#P_{n}-l$. This concludes Claim 5.15(3), (4) and the entire proof. ∎
@@
(5.3) (#$\operatorname{\mathbb{P}}\Big{(}\#P_{n+1}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})<k-j\,\Big{|}\,\#P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})=k\Big{)}\leq(4/\sqrt{N})^{j+1}.$

### v2

@@ 4. Schottky sets and random walks
$\zeta(S)$ $=\sup\big{\{}\zeta(I_{1}),\ldots,\zeta(I_{N}),\zeta(J_{1}),\ldots,\zeta(J_{N}),\big{\}}$
@@
$\big{(}\mathcal{A}:=\{i:\textrm{$I_{i}$ contains $x$ or $y$}\}\big{)}\leq 2.$
@@
Hence, $\mathcal{A}\cup\mathcal{B}$ has at most $\sqrt{N}+2<N$ elements, and we can pick an index $i$ outside it. For that $i$, we conclude that $d(gx,gy)<\operatorname{Leb}(J_{i})\leq 1/\sqrt{N}$ for each $g\in\mathfrak{S}(I_{i},J_{i})$. Since $G$ intersects with $\mathfrak{S}(I_{i},J_{i})$, we conclude that $\inf_{g\in G}d(gx,gy)<1/\sqrt{N}$. By sending $N$ to infinity, we conclude that the action of $G$ is proximal. ∎
@@ 4.1. Exponential Synchronization
First, let us write $S=\mathfrak{S}(I_{1},J_{1})\cup\ldots\cup\mathfrak{S}(I_{N},J_{N})$ for some essentially disjoint sets $I_{1},\ldots,I_{N},J_{1},\ldots,J_{N}$. Recall that elements of $\mathfrak{S}(I_{i},J_{i})$ send $\mathcal{I}$ into $J_{i}$. $(\ast)$ Note that $wJ_{1},\ldots,w_{J_{N}}$ are disjointly contained in $w\mathcal{I}$. Hence, the sum of their “size” is no greater than that of $w\mathcal{I}$, which implies that
@@
Now fixing $0\leq k\leq n-1$ and the choices of $\{s_{i}:1\leq i\leq k\}$, we observe that
@@
$\operatorname{\mathbb{P}}\big{(}\{x,y\}\cap w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\mathcal{I}=\emptyset\big{)}\geq 1-e^{-n}$
@@
holds regardless of the choices of $s_{i}$’s. Furthermore, when $0\leq k\leq n-1$ and $\{s_{i}:1\leq i\leq k\}$ are given,
@@
$\operatorname{\mathbb{P}}\big{(}\{x,y\}\cap W_{n}\mathcal{I}=\emptyset\big{)}\geq\operatorname{\mathbb{P}}\left(E_{1}\cup\ldots\cup E_{n}\right)\geq 1-(2/N)^{n}\geq 1-e^{-n}.$
@@ 4.2. Probabilistic Tits alternative
$\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$ intersect}\big{\}}\leq 3N\sqrt{N}.$
@@
$\geq\#\big{\{}(I_{i},gJ_{j}):I_{i}\cap gJ_{j}\neq\emptyset,I_{i}\in\mathcal{C}\big{\}}$
$\geq\#\mathcal{C}\cdot\min_{I_{i}\in\mathcal{C}}\#\{J_{j}:I_{i}\cap gJ_{j}\neq\emptyset\}$
$\geq\#\mathcal{C}\sqrt{N}$
@@
$\#\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$ intersect}\big{\}}$ $\leq\#\mathcal{C}\cdot N+(N-\#\mathcal{C})\cdot\sqrt{N}$
@@
$\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$ intersect}\big{\}}\leq 3\zeta^{2}N\sqrt{N}.$
@@
Hence, if $i\notin\mathcal{C}^{(l,m)}$ for each $(l,m)\in\{1,\ldots,\zeta\}^{2}$, then $\{j:I_{i}\cap gJ_{j}\neq\emptyset\}$ has cardinality $\leq\zeta^{2}\sqrt{N}$. Since $\mathcal{C}^{(l,m)}$ has cardinality at most $2\sqrt{N}$ for each $(l,m)$, we conclude that $\mathcal{C}$ consists of at most $2\zeta^{2}\sqrt{N}$ sets among $\{I_{1},\ldots,I_{N}\}$.
@@
$\#\big{\{}(i,j)\in\{1,\ldots,N\}^{2}:\textrm{$I_{i}$ and $gJ_{j}$ intersect}\big{\}}$ $\leq\#\mathcal{C}\cdot N+(N-\#\mathcal{C})\cdot\zeta^{2}\sqrt{N}$
@@
[…]
@@
$E_{1}$ $=\big{\{}s_{n+1}w_{n+1}\cdots s_{2n}w_{2n}\cdot w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\bar{\mathcal{I}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}\big{\}},$
$E_{2}$ $=\big{\{}t_{n+1}v_{n+1}\cdots t_{2n}v_{2n}\cdot v_{0}t_{1}v_{1}\cdots t_{n}v_{n}\bar{\mathcal{I}^{\prime}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}^{\prime}\big{\}},$
$E_{3}$ $=\big{\{}s_{n+1}w_{n+1}\cdots s_{2n}w_{2n}\cdot v_{0}t_{1}v_{1}\cdots t_{n}v_{n}\bar{\mathcal{I}^{\prime}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}\big{\}},$
$E_{4}$ $=\big{\{}t_{n+1}v_{n+1}\cdots t_{2n}v_{2n}\cdot w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\bar{\mathcal{I}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}^{\prime}\big{\}},$
$E_{5}$ $=\big{\{}s_{n+1}w_{n+1}\cdots s_{2n}w_{2n}\cdot v_{2n}^{-1}t_{2n}^{-1}\cdots v_{n+1}^{-1}t_{n+1}^{-1}\cdot\overline{S^{1}\setminus\mathcal{I}^{\prime}}\subseteq\operatorname{\operatorname{int}}\mathcal{I}\big{\}},$
$E_{6}$ $=\big{\{}v_{n}^{-1}t_{n}^{-1}\cdots v_{1}^{-1}t_{1}^{-1}v_{0}^{-1}\cdot w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\cdot\bar{\mathcal{I}}\subseteq\operatorname{\operatorname{int}}(S^{1}\setminus\mathcal{I}^{\prime})\big{\}}.$
@@
$I^{(1)}$ $=(s_{n+1}w_{n+1}\cdots s_{2n}w_{2n})^{-1}(S^{1}\setminus\operatorname{\operatorname{int}}\mathcal{I}),$
$I^{(2)}$ $=(t_{n+1}v_{n+1}\cdots t_{2n}v_{2n})^{-1}\cdot(S^{1}\setminus\operatorname{\operatorname{int}}\mathcal{I}^{\prime}),$
$J^{(1)}$ $=w_{0}s_{1}w_{1}\cdots s_{n}w_{n}\overline{\mathcal{I}},$
$J^{(2)}$ $=v_{0}t_{1}v_{1}\cdots t_{n}v_{n}\overline{\mathcal{I}^{\prime}}.$
@@
[…]
@@ 4.3. Local contraction
$\cdots\times\Omega_{3}\times\Omega_{2}\times\Omega_{1}=:\Omega,$
@@ 5. Pivoting technique
It suffices to prove this for $n$ being a multiple of $3m$. Indeed, for $n=3mk+l$ ($1\leq l\leq 3m-1$) we can treat as follows: we first take $\Omega_{3mk}$, $\mathcal{P}_{mk}$, $(w_{i})_{i},(r_{i},s_{i},t_{i})_{i}$ using the proposition and consider $(G^{l},\mu^{l})$ (where $G=\operatorname{Homeo}(S^{1})$). And then we define
$\Omega_{3mk+l}$ $=\Omega_{mk}\times G^{l},$
$\mathcal{P}_{3mk+l}$ $=\mathcal{P}_{mk}\times G^{l}=\{\mathcal{E}_{\alpha}\times(g_{1},\ldots,g_{l}):\mathcal{E}_{\alpha}\in\mathcal{P}_{3mk},(g_{1},\ldots,g_{l})\in G^{l}\}$
@@
$w_{0}$ $=\prod_{i=1}^{i(1)-1}g_{i}=\nu_{1}\cdots\nu_{i(1)-1},$
[…]
$w_{\lfloor\epsilon n/10m\rfloor}$ $=\prod_{i=i(\lfloor\epsilon n/10m\rfloor)+1}^{n/3m}g_{i}=\nu_{i(\lfloor\epsilon n/10m\rfloor)+1}\cdots\nu_{n/3m}$
@@
First, note that the elements of $P_{u}$ smaller than or equal to m1m-1 must have been acquired no later than step m1m-1, and then must have never been lost thereafter. Hence, they all belong to $P_{m-1}$. Meanwhile, all elements $P_{m-1}$ should have remained till step $u$ for the following reason. If an element of $P_{m-1}\subseteq\{1,\ldots,m-1\}$ was lost at some step $n\in\{m,m+1,\ldots,u\}$, it would mean that scenario (2-B) was the case at step $n$, with $k=\max\mathcal{Q}$ being smaller than m1m-1. This means that $P_{n}$ lost not only $P_{m-1}$ but also $m$, which contradicts $P_{u}\ni m$. Hence the claim follows.
@@
    l=m1l=m-1: this means that scenario (2-A) was the case at both step $l$ and step $m=l+1$. Hence, $J(r_{l+1})\subseteq W_{l}^{-1}L_{l}:=w_{l}^{-1}\big{(}S^{1}\setminus I(t_{l})\big{)}$ must hold. This implies
@@
    l<m1l<m-1: in this case, $P_{m-1}=P_{u}\cap\{1,\ldots,m-1\}\subseteq\{1,\ldots,l\}$ does not contain m1m-1 so scenario (2-B) must have been the case at step n=m1n=m-1. Still, $P_{m-1}=P_{u}\cap\{1,\ldots,m-1\}$ contains an element $l$ so scenario (2-B-ii) is ruled out. Thus, scenario (2-B-i) was the case and $l$ must have been the maximum element of $\mathcal{Q}$. This leads to $L_{m-1}:=W_{l-1}r_{l}s_{l}\mathcal{I}$. We now know that scenario (2-A) was the case at step $n=m$, which implies $J(r_{m})\subseteq W_{m}^{-1}L_{m-1}$. Hence, we conclude
@@
Let $S=\cup_{i=1}^{N}\mathfrak{S}(I_{i},J_{i})$ for essentially disjoint subsets $\{I_{i},J_{i}\}_{i}$. Note that the set $P_{n-1}$ and the interval $L_{n-1}$ are determined from the fixed inputs. Now at step n1n-1 of the pivotal set construction, three possibilities arise:
@@
Based on our estimates for the probabilities for $I(t_{n})\notin\mathcal{C}(w_{n};S)$ and $J(r_{n})\subseteq W_{n-1}^{-1}L_{n-1}$ in the above three cases, we can conclude that $P_{n+1}=\#P_{n}+1$ happens for probability at least $1-4\zeta^{2}/\sqrt{N}$. ∎
@@
$\textrm{$s_{i}=\bar{s}_{i}$ for each $i\in\{1,\ldots,n\}\setminus P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})$},$
@@
As the base case, we have observed Item (4) for $l=k$. For general $k<l\leq n$, we will start by assuming Item(4) for l1l-1. Recall the conditions for scenario (2-A) at step $l$, for the input $\mathbf{s}$:
@@
[…]
@@
$\operatorname{\mathbb{P}}_{\mathcal{E}}\big{(}B_{l}\big{)}$ $=\int_{(\mathbf{r},\mathbf{s},\mathbf{t})\in B_{l-1}}\operatorname{\mathbb{P}}_{s_{i(l)}\sim\mu}\big{(}s_{i(l)}\notin A_{l-1}\,\big{|}\,s_{i(l-1)},\ldots,s_{i(1)},r_{n+1},t_{n+1}\big{)}\,d\mu(s_{i(l-1)})\cdots d\mu(s_{i(1)})\,d\mu(r_{n+1})\,d\mu(t_{n+1})$
@@
It remains to prove the claim. The claim regarding $A_{0}$ was already established in Lemma 5.10. That means, regardless of the values of $(s_{i(M)},\ldots,s_{i(1)})$, we proved that the probability for $(r_{n+1},t_{n+1})$ to satisfy $P_{n+1}=\#P_{n}+1$ is at least $1-4\zeta^{2}/\sqrt{N}$. We will prove something more: we claim that the candidates for $r_{n+1},t_{n+1}$ that make $P_{n+1}=\#P_{n}+1$ are independent of $(s_{i(M)},\ldots,s_{i(1)})$. When restricted to $\mathcal{E}$, $P_{n+1}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})=\#P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})+1$ holds if and only if $I(t_{n+1})\in\mathcal{C}(w_{n+1};S)$ and $J(r_{n+1})\subseteq W_{n}^{-1}L_{n}$. Here,
@@
are fixed throughout $\mathcal{E}$. This is why $P_{n+1}=\#P_{n}+1$ depends on the choice of $r_{n+1}$ and $t_{n+1}$, regardless of the values of $s_{i(1)},\ldots,s_{i(M)}$. This settles Claim 5.15(1), (2) and also the construction of $A_{0}$.
@@
$A_{l}$ $=\Big{\{}s\in S:I(s)\notin\mathcal{C}\big{(}t_{i(l)}w_{i(l)}\cdot W_{i(l)}^{-1}W_{n+1};S\big{)}\Big{\}}$
$=\Big{\{}s\in S:I(s)\notin\mathcal{C}\big{(}t_{i(l)}w_{i(l)}\cdot(r_{i(l)+1}s_{i(l)+1}t_{i(l)+1}w_{i(l)+1})\cdots(r_{n}s_{n}t_{n}w_{n})\cdot(r_{n+1}s_{n+1}t_{n+1}w_{n+1});S\big{)}\Big{\}}.$
@@
(5.2) $\big{\{}j:\bar{I}(s_{i(l)})\cap t_{i(l)}w_{i(l)}\cdot W_{i(l)}^{-1}W_{n+1}\bar{J}_{j}\neq\emptyset\big{\}}\leq\zeta^{2}\sqrt{N}.$
@@
$\big{\{}j:\bar{I}(t_{i(l+1)})\cap(W_{i(l+1)}w_{i(l+1)}^{-1})^{-1}\cdot(W_{i(l)-1}r_{i(l)}s_{i(l)})\cdot t_{i(l)}w_{i(l)}\cdot W_{i(l)}^{-1}W_{n+1}\bar{J}_{j}\neq\emptyset\big{\}}\leq\sqrt{N}.$
In other words, $I(t_{i(l+1)})\notin\mathcal{C}\big{(}w_{i(l+1)}\cdot(W_{i(l+1)})^{-1}\cdot W_{n+1};S\big{)}$ holds true. This implies that the set $\mathcal{Q}$ in scenario (2-B) at step $n+1$ contains $i(l+1)$. Hence, $P_{n+1}(\mathbf{r},\mathbf{s},\mathbf{t})$ contains $P_{n}(\mathcal{E})\cap\{1,\ldots,i(l+1)\}=\{i(M)<\ldots<i(l+1)\}$ at least, which leads to the inequality $P_{n+1}\geq\#P_{n}-l$. This concludes Claim 5.15(3), (4) and the entire proof. ∎
@@
(5.3) $\operatorname{\mathbb{P}}\Big{(}\#P_{n+1}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})<k-j\,\Big{|}\,\#P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})=k\Big{)}\leq(4/\sqrt{N})^{j+1}.$